• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Rvachev

    Log entry time 08:29:39 on December 17,1999

    Entry number 31301

    keyword=Shift summary

    12:00 Waiting for beam
    1:10 Got beam
    1:30 Bob came in and is doing some tests.
    2:40 Problem with trigger partially fixed (at least we don't see it now). It
    turns out it is safer to repeat elastic normalization points, since it is not clear
    how to correct the data.
    3:05 Start normalization runs.

    moving to Q2=1.31 elastic kinematics.
    Cycling of EQ2 and EQ3 to 1600A done
    Hadron arm is a lumonisity monitor

    For helium raster size is 5X4 at MCC, fan is at 20%

    1377 12C Q2=1.31 FF on 5uA
    1378 He3 Q2=1.31 FF off current not very stable 1.2 uA
    1379 He3 Q2=1.31 FF off current not too stable 1.2uA
    1380 He3 Q2=1.31 FF off current more stable 1.2 uA
    1381 He3 Q2=1.90 FF off 1.5uA; short run
    1382 He3 Q2=1.90 FF off 1.5uA
    1383 He3 Q2=1.90 FF off 1.5uA (while waiting for Hall B finish normalization run)
    From now on current is pretty stable
    1384 He3 Q2=1.90
    1385 12C Q2=1.90 FF on
    1386 12C Q2=1.90 (hadron only;e-arm moving) FF on
    1387 junk
    1388 12C Q2=5.01 FF on
    1389 He3 Q2=5.01 FF off 1.5uA
    1390 He3 Q2=5.01 FF off 1.5uA
    1391 He3 Q2=15.00 FF on 15uA (MCC refuses to give more! Arun was called)

    6:50 MCC got a serious problem with a cavity. They have to take it out of the tune (takes time)
    7:00 We did not include the 1.0055 fudge factor in the electron momentum Marty
    gave us yesterday at Q2=15. We recycle EQ2 and EQ3 to 1600A.

    7:50 Got beam
    8:00 Maximum current MCC is able to deliver is 45uA, but we could go to ~70uA.
    For now we accept 45uA, and we noticed a problem with T1 and T2, and leave it to
    the next shift crew to investigate.
    1392 He3 Q2=15.00 45uA (test run)
    1393 He3 Q2=15.00 45uA (test run)